# Anchors Aweigh! Sail for Optimal Unified Multi-Modal Representations

### v1

@@ Abstract
Multimodal learning plays a crucial role in enabling machine learning models to fuse and utilize diverse data sources, such as text, images, and audio, to support a variety of downstream tasks. A unified representation across various modalities is particularly important for improving efficiency and performance. Recent binding methods, such as ImageBind (Girdhar et al., 2023), typically use a fixed anchor modality to align multimodal data in the anchor modal embedding space. In this paper, we mathematically analyze the fixed anchor binding methods and uncover notable limitations: (1) over-reliance on the choice of the anchor modality, (2) failure to capture intra-modal information, and (3) failure to account for inter-modal correlation among non-anchored modalities. To address these limitations, we propose CentroBind, a simple yet powerful approach that eliminates the need for a fixed anchor; instead, it employs dynamically adjustable centroid-based anchors generated from all available modalities, resulting in a balanced and rich representation space. We theoretically demonstrate that our method captures three crucial properties of multimodal learning: intra-modal learning, inter-modal learning, and multimodal alignment, while also constructing a robust unified representation across all modalities. Our experiments on both synthetic and real-world datasets demonstrate the superiority of the proposed method, showing that dynamic anchor methods outperform all fixed anchor binding methods as the former captures more nuanced multimodal interactions.
@@ 1 Introduction
Multimodal alignment is defined as identifying and exploiting relationships and correspondences between sub-components of instances from multiple modalities (e.g., text, image, audio) to establish meaningful connections between their representations (Baltrušaitis et al., 2018). This process allows machine learning models to analyze heterogeneous data holistically, facilitating comprehensive decision-making. A common approach is learning a shared embedding space (Tu et al., 2022; Girdhar et al., 2023; Liang et al., 2024b; Zhu et al., 2024), which aims to project data from multiple modalities into a common embedding space by clustering similar items together for direct comparison and linkage. This approach leverages well-trained single-modal embeddings, aligning them with auxiliary objective functions like contrastive (Oord et al., 2018) or triplet loss (Wang et al., 2020) to minimize distances between similar items and maximize distances between dissimilar ones across modalities.
Instead of training separate models for each modality, ImageBind (Girdhar et al., 2023) pairs images with other modalities and projects them into the common image embedding space. Similarly, Zhu et al. (2024) shows that pairing texts with other modalities (LanguageBind) improves retrieval task performance when language is specified as the anchored modality. This approach has inspired various “-Bind” methods tailored to align different modalities for specific domains, such as molecular modeling (Xiao et al., 2024), medical imaging (Gao et al., 2024), brain signals (Yang et al., 2024b), and music selection for videos (Teng et al., 2024). These models commonly use image or text as the anchor embedding due to the abundance of data, with other modalities projected into this anchor representation.
We can define the aforementioned approaches as Fixed-Anchor-Bind (FABind) method, where the embedding space of the primary anchor modality remains fixed during the alignment process. Generally, the “-Bind”-like approaches maximize mutual information $I({\bf Z}_{1};{\bf Z}_{i})$ between the representation ${\bf Z}_{1}$ of the anchor modality and the representations ${\bf Z}_{i},i\in\{2,\cdots,M\}$ of other modalities. Although these approaches are practically useful and widely adopted in learning unified multimodal representation, we theoretically and empirically demonstrate the limitations in FABind .
@@
First, selecting an anchor modality is crucial but challenging, as it depends on both embedding quality and task suitability. Common choices like images or text may still be suboptimal, especially for less common modalities lacking high-quality embeddings. Second, fixing an anchor can result in lost semantic information that might be better represented by other modalities. For instance, while text may describe ’a dog barks loudly,’ sound could reveal mood, and an image could add facial expression, which fixed alignment might miss. Third, optimizing only for anchor-to-other-modality overlooks information similarity between non-anchored modalities, leading to a loss of complementary insights. Addressing this among these non-anchored modalities would disrupt overall multimodal alignment, contradicting the primary goal of CentroBind. We formally analyze deficiencies of the fixed anchor bind approaches in Section 2.
(a) FABind vs. CentroBind.
(b) Graphical illustration of CentroBind.
@@
Figure 1: Fixed anchor bind methods (FABind) binds representations to the fixed anchor modality, while CentroBind uses dynamic anchors. $\rm(a)$ Colors and shapes represent different modalities and semantic information, respectively. $\mathcal{Z}$ denotes the unified representation space. $\rm(b)$ CentroBind forms dynamic anchors from the centroids of positive augmentation pairs.
Dynamic anchor alignment.
@@
We propose a simple yet effective modification by replacing fixed anchors with “dynamic” centroid-based anchors computed from paired samples. Our method, CentroBind, described in Section 3 removes the need for selecting a fixed anchor modality, instead calculates the centroid of all modality representations and generates an anchor representation, as shown in Figure 1(a). Encoders are then trained to minimize the ensemble of InfoNCE loss (Oord et al., 2018) between this dynamic anchor and other representations, using the centroid as the anchor. The main intuition is that a desirable anchor should be representative of all modalities, capturing the most comprehensive information, with well-trained encoders producing representations that naturally cluster around this shared centroid, reflecting their underlying semantic alignment.
@@
Our theoretical analysis demonstrates that CentroBind effectively addresses three critical components of multimodal learning: 1) intra-modal mutual information, 2) inter-modal mutual information, and 3) multimodal alignment via embedding similarity. By incorporating these elements within the ‘anchor alignment’ framework, CentroBind benefits from the significant advantages of dynamic anchors across both synthetic and real-world datasets in retrieval and classification tasks. In a sense, our approach yields an ideal unified representation space, supported by the perspective of Huh et al. (2024), which conjectures that multimodal representations align as they move toward a platonic representation—an ideal form that captures the semantic information of all modalities simultaneously.
@@ 2 Problem formulation
In this section, we describe general representation learning and representation binding problems in multimodal learning. Then, we analyze fixed-anchor-bind (FABind) methods such as ImageBind (Girdhar et al., 2023), that bind multimodal representation to a fixed modality of choice.
@@ 2.1 Representation learning framework
Given $M$ datasets $\mathcal{D}=\{\mathcal{D}_{i}\}_{i=1}^{M}$, let $\mathcal{D}_{i}=\{({\boldsymbol{x}}_{i,j},{\boldsymbol{y}}_{i,j})\}_{j=1}^{N_{i}}$ be the dataset from the $i$-th modality, where ${\boldsymbol{x}}_{i,j}\in\mathcal{X}_{i}$ and ${\boldsymbol{y}}_{i,j}\in\mathcal{Y}_{i}$ are respectively the $j$-th input instance (e.g., feature vector) and the corresponding label in $i$-th modality, and we assume that $({\boldsymbol{x}}_{i,j},{\boldsymbol{y}}_{i,j})\overset{\rm i.i.d.}{\sim}P_{{\bf X}_{i},{\bf Y}_{i}}$. [^1] In self-supervised learning, labels might not exist, which corresponds to the case that ${\boldsymbol{y}}_{i,j}$’s are empty. We assume that $j$ indexes paired samples among modalities. For instance, ${\boldsymbol{x}}_{1,c}$ and ${\boldsymbol{x}}_{2,c}$ are features having similar semantic information (e.g., dog image and dog sound) in $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$. The goal of representation learning is to build $M$ encoders $f_{i}:\mathcal{X}_{i}\to\mathcal{Z}_{i}$ for each modality, which maps the input instances ${\boldsymbol{x}}_{i,j}$ to its embedding ${\boldsymbol{z}}_{i,j}=f_{i}({\boldsymbol{x}}_{i,j})$, preserving as much information about ${\boldsymbol{x}}_{i,j}$ as possible.
@@
For the uni-modal case ($M=1$), keeping maximum information about ${\boldsymbol{x}}_{1,j}$ at its embedding ${\boldsymbol{z}}_{1,j}$ is generally preferred based on the “InfoMax” principle (Linsker, 1988), under which the objective is to maximize mutual information $I({\bf X}_{i};f({\bf X}_{i}))$ between ${\bf X}_{i}$ and $f({\bf X}_{i})$. Throughout the paper, we call $I({\bf X}_{i};f({\bf X}_{i}))$ intra information on ${\bf X}_{i}$. For the multimodal case ($M\geq 2$), on top of the InfoMax principle, “minimal sufficiency” is proposed in (Tian et al., 2020), which suggests maximizing shared information $I(f_{i}({\bf X}_{i});f_{l}({\bf X}_{l}))$ between $f_{i}({\bf X}_{i})$ and $f_{l}({\bf X}_{l})$, while minimizing the unique information $I({\bf X}_{i};f_{i}({\bf X}_{i})|\{{\bf X}_{l}\}_{l\neq i})$. Although minimal sufficiency often leads to an efficient encoder with a better performance in numerous multimodal downstream tasks, it is not always a good strategy as there exist exceptions where the unique information on an individual modality is crucial (Liang et al., 2024b; Wang et al., 2022). In other words, the optimality of minimal sufficiency is task-dependent. To avoid task dependency, we do not consider minimal sufficiency; instead, we maximize intra and shared information without reducing unique information. Next, we formalize the notion of sufficient embedding.
Definition 1 ($\mathcal{Z}_{i}$-Sufficient embedding of ${\bf X}_{i}$ for ${\bf X}_{l}$).
@@
For embedding space $\mathcal{Z}$, the embedding $f_{i}({\bf X}_{i})$ is $\mathcal{Z}_{i}$-sufficient for ${\bf X}_{l}$ if and only if
$f_{i}\in{\hbox{arg}}\max_{f:\mathcal{X}_{i}\to\mathcal{Z}_{i}}I(f({\bf X}_{i});{\bf X}_{l}),$ (1)
and we call $f_{i}$ sufficient encoder of ${\bf X}_{i}$ for ${\bf X}_{l}$.
@@
We note that if $i=l$, the sufficient encoder provides embeddings with maximum intra information, and if $i\neq l$, it gives embeddings with maximum shared information between $i$-th and $l$-th modalities. [^2] With a proper choice of $\mathcal{Z}_{i}$ ensuring $\max_{f:\mathcal{X}_{i}\to\mathcal{Z}_{i}}I(f({\bf X}_{i});{\bf X}_{l})=I({\bf X}_{i};{\bf X}_{l})$, Definition 1 says that ${\boldsymbol{z}}_{i,j}=f_{i}({\boldsymbol{x}}_{i,j})$ is a sufficient statistic (Polyanskiy & Wu, 2024) of ${\boldsymbol{x}}_{i,j}$ for ${\boldsymbol{x}}_{l,j}$ as the encoding entails no information loss.
@@
In the context of contrastive representation learning, with a goal of attaining sufficient encoders in Definition 1, InfoNCE loss $I_{\rm NCE}(X;Y)$ is often employed since it relates to mutual information. Specifically, InfoNCE provides a lower bound on mutual information, i.e., $I({\bf X};{\bf Y})\geq-I_{\rm NCE}({\bf X};{\bf Y})$ (Oord et al., 2018), thus minimizing InfoNCE leads to an increase in mutual information. InfoNCE loss between embeddings ${\bf U}$ and ${\bf V}$ can be written as follows:
$I_{\rm NCE}({\bf U};{\bf V}|\tau)$ $=\mathbb{E}_{P_{{\bf U},{\bf V}},\prod_{i=1}^{N}P_{{\bf V}_{i}}}\left[-\log\frac{\exp({\bf U}^{\top}{\bf V}/\tau)}{\exp({\bf U}^{\top}{\bf V}/\tau)+\sum_{i=1}^{N}\exp({\bf U}^{\top}{\bf V}_{i}/\tau)}\right],$ (2)
@@
where the expectation is taken with respect to $P_{{\bf U},{\bf V}}\prod_{i=1}^{N}P_{{\bf V}_{i}}$. Here, we say $({\bf U},{\bf V})\sim P_{{\bf U},{\bf V}}$ a positive pair and $({\bf U},{\bf V}_{i})\sim P_{\bf U}P_{{\bf V}_{i}}$ a negative pair. Moreover, $N\geq 1$ and $\tau>0$ are hyper-parameters, specifying the number of negative samples and the temperature parameter. For simplicity, we assume that embeddings are normalized (Wang & Isola, 2020) and are of the same dimensionality in this paper. Under the assumption of $\tau=1$, the exponent ${\bf U}^{\top}{\bf V}/\tau$ in (2) is the cosine similarity score between ${\bf U}$ and ${\bf V}$.
@@ 2.2 Binding representation spaces
In addition to the objective of intra and shared information, multimodal learning often takes into account multimodal alignment (Radford et al., 2021; Duan et al., 2022). Without multimodal alignment, each modality has its own embedding structure depending on its encoder. For example, embeddings of cat and dog images, respectively, locate around $(1,0)$ and $(0,2)$ in $\mathbb{R}^{2}$, whereas embeddings of cat and dog text can lie around $(0,2)$ and $(1,0)$. Such a misalignment can happen even for sufficient encoders (Definition 1), since the mutual information is invariant to one-to-one mappings (Polyanskiy & Wu, 2024).
To align multimodal embedding spaces, a unified representation space (Radford et al., 2021; Zhou et al., 2023) or multimodal alignment (Wang et al., 2023; Liang et al., 2024c) have gained attention and become important property in multimodal representation learning, in which embeddings of multimodal features having similar semantic should locate nearby. Among various techniques, efficient FABind methods have been proposed (see Appendix A for a summary of FABind methods) , such as ImageBind (Girdhar et al., 2023). Their main idea is to set the image modality as a fixed anchor modality, and they minimize the InfoNCE loss between the embeddings of the anchor modality and the other modalities. More broadly, FABind (e.g., ImageBind) aims to find encoders $f_{i}^{\rm FB}$ for all modalities except the anchor modality such that
$f^{\rm FB}_{i}={\hbox{arg}}\min_{f_{i}:\mathcal{X}_{i}\to\mathcal{Z}_{i}}I_{\rm NCE}(f_{1}({\bf X}_{1});f_{i}({\bf X}_{i})),~{}\forall i\in\{2,\cdots,M\},$ (3)
where $f_{1}$ is the encoder for the anchor modality (e.g., image encoder in ImageBind). Note that FABind freezes $f_{1}$, which is initialized by an existing pretrained model, during the optimization.

### v2

@@ Abstract
A unified representation space in multi-modal learning is essential for effectively integrating diverse data sources, such as text, images, and audio, to enhance efficiency and performance across various downstream tasks. Recent binding methods, such as ImageBind (Girdhar et al., 2023), typically rely on a single, fixed anchor modality for aligning multi-modal data. We mathematically analyze these fixed anchor binding methods and uncover significant limitations: (1) over-reliance on the choice of the anchor modality, (2) inadequate capture of intra-modal information, and (3) failure to account for cross-modal correlation among non-anchored modalities. To address these issues, we propose the need for adaptive anchor binding methods, exemplified by our framework CentroBind. The proposed method uses adaptively adjustable centroid-based anchors generated from all available modalities, leading to a balanced and rich representation space. We theoretically demonstrate that our approach captures three critical properties of multi-modal learning—intra-modal learning, inter-modal learning, and multi-modal alignment—while constructing a unified representation that spans all modalities. Experiments on both synthetic and real-world datasets show that adaptive anchor methods such as CentroBind consistently outperform fixed anchor binding methods, verifying our analysis.
Machine Learning, ICML
@@ 1 Introduction
Multi-modal alignment is defined as identifying and exploiting relationships and correspondences between multiple modalities (e.g., text, image, audio) viewing common phenomena to establish meaningful connections between their representations (Baltrušaitis et al., 2018). A common approach is learning a shared embedding space (Tu et al., 2022; Girdhar et al., 2023; Liang et al., 2024b; Zhu et al., 2024), which aims to project data from multiple modalities into a common embedding space by clustering similar items together for direct comparison and linkage. This approach leverages well-trained single-modal embeddings, aligning them with auxiliary objective functions like contrastive loss (Oord et al., 2018) or triplet loss (Schroff et al., 2015) to minimize distances between similar items and maximize distances between dissimilar ones across modalities.
Instead of training separate models for each modality, ImageBind (Girdhar et al., 2023) pairs images with other modalities and projects them into a common image embedding space. Similarly, Zhu et al. (2024) shows that pairing texts with other modalities (LanguageBind) improves cross-modal retrieval performance when language serves as the anchor modality. This approach has inspired various “-Bind” methods tailored to align different modalities for specific domains, such as molecular modeling (Xiao et al., 2024), medical imaging (Gao et al., 2024), brain signals (Yang et al., 2024b), and music selection for videos (Teng et al., 2024). These models commonly use image or text as the anchor embedding due to the abundance of data, with other modalities projected into this anchor representation.
We define the aforementioned approaches as Fixed Anchor Bind (FABind) methods, where the embedding space of the primary anchor modality remains fixed during the alignment process. Many “-Bind” approaches maximize mutual information $I({\bf Z}_{1};{\bf Z}_{i})$ between the representation ${\bf Z}_{1}$ of the anchor modality and the representations ${\bf Z}_{i},i\in\{2,\cdots,M\}$ of other modalities. Although these approaches are practically useful and widely adopted in learning unified multi-modal representation, they have limitations, as demonstrated through theory and experiment in this paper.
@@
First, selecting which modality to serve as an anchor is crucial but challenging, as it depends on both embedding quality and task suitability. Common choices like images or text can be suboptimal, especially when no single modality has the dominant information about the source. Second, fixing an anchor can result in loss of semantic information that is well represented only in other modalities. For instance, while text may describe ‘a dog barks loudly,’ sound could reveal mood, and an image could add facial expression, a holistic combination that a fixed alignment might miss. Third, optimizing only for anchor-to-other-modality overlooks information complementarity between non-anchored modalities. These challenges serve as the primary motivation for introducing adaptive anchor alignment as an alternative to FABind. We formally analyze deficiencies of FABind in Section 2.
@@
Figure 1: Fixed anchor binding methods (FABind) rely on a fixed anchor modality, while adaptive anchors (CentroBind) dynamically generate anchoring points from multi-modal batch samples. Colors and shapes represent modalities and semantic information, respectively, with $\mathcal{Z}$ as the unified representation space.
Figure 2: Graphical illustration of adaptive anchor alignment: Adaptive anchors are dynamically generated from current embeddings using an aggregation method (CentroBind) for every batch.
Adaptive anchor alignment.
We propose an alternative to fixed anchor alignment by replacing fixed anchors with “adaptive” anchors computed from paired multi-modal samples. Our proposed method, CentroBind, an example of adaptive anchor bind methods, described in Section 3, removes the need for selecting a fixed anchor modality, instead calculates the centroid over the aggregate of all modality’s representations and generates a multi-modal anchor representation, as shown in Figure 1. We note that various aggregation methods can be used for these embeddings, such as computing a weighted average when the relative importance or quality of the backbone models is known, or even learning the weight coefficients dynamically during training. Some of these options are explored, with experimental results provided in Appendix C.1.
@@
Encoders are then trained to minimize the ensemble of InfoNCE loss (Oord et al., 2018) between the representation of this adaptive anchor and the fixed modality representations. The main intuition is that a desirable anchor should be representative of all modalities, capturing the most comprehensive information, with well-trained encoders producing representations that naturally cluster around this shared centroid, reflecting their underlying semantic alignment.
@@
Our theoretical analysis demonstrates that CentroBind effectively addresses three critical components of multi-modal learning: 1) capturing intra-modal mutual information, 2) learning inter-modal mutual information, and 3) performing multi-modal alignment by maximizing embedding similarity measures. By incorporating these elements, CentroBind outperforms other multi-modal alignment methods, as shown empirically on both synthetic and real-world datasets in retrieval and classification tasks. The proposed approach yields a unified representation space and, in the perspective of (Huh et al., 2024), who contends that multi-modal representations better align as they move toward a platonic representation that captures the semantic information of all modalities simultaneously.
@@ 2 Problem Formulation
In this section, we describe general representation learning and binding problems in multi-modal learning. Then, we analyze fixed-anchor-bind (FABind) methods such as ImageBind (Girdhar et al., 2023), that bind multi-modal representations to a user-selected fixed modality.
@@ 2.1 Representation learning framework
Given $M$ datasets $\mathcal{D}=\{\mathcal{D}_{i}\}_{i=1}^{M}$, let $\mathcal{D}_{i}=\{({\boldsymbol{x}}_{i,j},{\boldsymbol{y}}_{i,j})\}_{j=1}^{N_{i}}$ be the dataset from the $i$-th modality, where ${\boldsymbol{x}}_{i,j}\in\mathcal{X}_{i}$ and ${\boldsymbol{y}}_{i,j}\in\mathcal{Y}_{i}$ are respectively the $j$-th input instance (e.g., feature vector) and the corresponding label in $i$-th modality, and we assume that $({\boldsymbol{x}}_{i,j},{\boldsymbol{y}}_{i,j})\overset{\rm i.i.d.}{\sim}P_{{\bf X}_{i},{\bf Y}_{i}}$. [^1] In self-supervised learning, labels might not exist, which corresponds to the case that ${\boldsymbol{y}}_{i,j}$ are empty. We assume that $j$ indexes paired samples among modalities. For instance, ${\boldsymbol{x}}_{1,c}$ and ${\boldsymbol{x}}_{2,c}$ are features having similar semantic information (e.g., dog image and dog sound) in $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$. The goal of representation learning is to build $M$ encoders $f_{i}:\mathcal{X}_{i}\to\mathcal{Z}_{i}$ for each modality, which maps the input instances ${\boldsymbol{x}}_{i,j}$ to its embedding ${\boldsymbol{z}}_{i,j}=f_{i}({\boldsymbol{x}}_{i,j})$, preserving as much information about ${\boldsymbol{x}}_{i,j}$ as possible.
@@
For the uni-modal case ($M=1$), keeping maximum information about ${\boldsymbol{x}}_{1,j}$ at its embedding ${\boldsymbol{z}}_{1,j}$ is generally preferred based on the “InfoMax” principle (Linsker, 1988), under which the objective is to maximize mutual information $I({\bf X}_{i};f({\bf X}_{i}))$ between ${\bf X}_{i}$ and $f({\bf X}_{i})$. Throughout the paper, we call $I({\bf X}_{i};f({\bf X}_{i}))$ intra information on ${\bf X}_{i}$. For the multi-modal case ($M\geq 2$), on top of the InfoMax principle, “minimal sufficiency” is proposed in (Tian et al., 2020b), which suggests maximizing shared information $I(f_{i}({\bf X}_{i});f_{l}({\bf X}_{l}))$ between $f_{i}({\bf X}_{i})$ and $f_{l}({\bf X}_{l})$, while minimizing the unique information $I({\bf X}_{i};f_{i}({\bf X}_{i})|\{{\bf X}_{l}\}_{l\neq i})$. Although minimal sufficiency often leads to an efficient encoder with better performance in numerous multi-modal downstream tasks, it is not always a good strategy as there exist exceptions where the unique information on an individual modality is crucial (Liang et al., 2024b; Wang et al., 2022a). In other words, the optimality of minimal sufficiency is task-dependent. To avoid task dependency, we do not consider minimal sufficiency; instead, we maximize intra and shared information without reducing unique information. Next, we formalize the notion of sufficient embedding.
Definition 2.1 ($\mathcal{Z}_{i}$-Sufficient embedding of ${\bf X}_{i}$ for ${\bf X}_{l}$).
@@
For an embedding space $\mathcal{Z}_{i}$, the embedding $f_{i}({\bf X}_{i})$ is $\mathcal{Z}_{i}$-sufficient for ${\bf X}_{l}$ if and only if the embedding $f_{i}({\bf X}_{i})$ achieves the maximum mutual information between $f_{i}({\bf X}_{i})$ and ${\bf X}_{i}$. Specifically,
$f_{i}\in{\hbox{arg}}\max_{f:\mathcal{X}_{i}\to\mathcal{Z}_{i}}I(f({\bf X}_{i});{\bf X}_{l}).$ (1)
We call $f_{i}$ sufficient encoder of ${\bf X}_{i}$ for ${\bf X}_{l}$.
@@
We note that if $i=l$, the sufficient encoder provides embeddings with maximum intra information, and if $i\neq l$, it gives embeddings with maximum shared information between $i$-th and $l$-th modalities. [^2] With a proper choice of $\mathcal{Z}_{i}$ ensuring $\max_{f:\mathcal{X}_{i}\to\mathcal{Z}_{i}}I(f({\bf X}_{i});{\bf X}_{l})=I({\bf X}_{i};{\bf X}_{l})$, Definition 2.1 says that $f_{i}({\boldsymbol{x}}_{i,j})$ is a sufficient statistic (Polyanskiy & Wu, 2024) of ${\boldsymbol{x}}_{i,j}$ for ${\boldsymbol{x}}_{l,j}$ as the encoding entails no information loss.
@@
In the context of contrastive representation learning having the goal of attaining sufficient encoders in Definition 2.1, InfoNCE loss $I_{\rm NCE}(X;Y)$ is often employed since it relates to mutual information. Specifically, InfoNCE provides a lower bound on mutual information, i.e., $I({\bf X};{\bf Y})\geq-I_{\rm NCE}({\bf X};{\bf Y})$ (Oord et al., 2018), and thus minimizing InfoNCE leads to an increase in mutual information. InfoNCE loss between embeddings ${\bf U}$ and ${\bf V}$ can be written as follows:
$I_{\rm NCE}({\bf U};{\bf V}|\tau)$ $=\mathbb{E}_{P}\left[-\log\frac{e^{\frac{{\bf U}^{\top}{\bf V}}{\tau}}}{e^{\frac{{\bf U}^{\top}{\bf V}}{\tau}}+\sum_{i=1}^{N}e^{\frac{{\bf U}^{\top}{\bf V}_{i}}{\tau}}}\right],$ (2)
@@
where the expectation is taken with respect to the distribution $P=P_{{\bf U},{\bf V}}\prod_{i=1}^{N}P_{{\bf V}_{i}}$. Here, we say $({\bf U},{\bf V})$ is a positive pair if $({\bf U},{\bf V})\sim P_{{\bf U},{\bf V}}$ and $({\bf U},{\bf V})$ is a negative pair if $({\bf U},{\bf V}_{i})\sim P_{\bf U}P_{{\bf V}_{i}}$. In (2), $N\geq 1$ and $\tau>0$ are hyper-parameters, specifying the number of negative samples and the temperature parameter. For simplicity, in this paper, we assume that embeddings are normalized (Wang & Isola, 2020) to unit vectors and are of the same dimensionality. Then, the exponent ${\bf U}^{\top}{\bf V}/\tau$ in (2) is proportional the cosine similarity score between ${\bf U}$ and ${\bf V}$.
@@ 2.2 Binding representation spaces
In addition to the objective of capturing intra and shared information, multi-modal learning often takes into account multi-modal alignment (Radford et al., 2021; Duan et al., 2022). Without multi-modal alignment, each modality can only access its own embedding structure depending on its encoder. For example, embeddings of cat and dog images, respectively, locate around $(1,0)$ and $(0,2)$ in $\mathbb{R}^{2}$, whereas embeddings of cat and dog text can lie around $(0,2)$ and $(1,0)$. Such a misalignment can happen even for sufficient encoders (Definition 2.1), since the mutual information is invariant to one-to-one mappings (Polyanskiy & Wu, 2024).
To align multi-modal embedding spaces, a unified representation space (Radford et al., 2021; Zhou et al., 2023) or multi-modal alignment (Wang et al., 2023; Liang et al., 2024c) have been proposed for multi-modal representation learning, in which embeddings of multi-modal features having similar semantic should near each other in embedding space. Several FABind methods have been proposed (see Appendix A for a summary of FABind methods) that include ImageBind (Girdhar et al., 2023). ImageBind sets the image modality as the fixed anchor modality, and then InfoNCE loss is minimized between the embeddings of the anchor modality and the other modalities. FABind (e.g., ImageBind) aims to find encoders $f_{i}^{\rm FB}$ for all modalities, except the anchor modality, such that for all $i\in\{2,\cdots,M\}$,
$f^{\rm FB}_{i}\in{\hbox{arg}}\min_{f_{i}:\mathcal{X}_{i}\to\mathcal{Z}_{i}}I_{\rm NCE}(f_{1}({\bf X}_{1});f_{i}({\bf X}_{i})),$ (3)
where $f_{1}$ is the encoder for the anchor modality (an image encoder in ImageBind). Note that FABind freezes $f_{1}$, initialized by an existing pretrained model, during the optimization.